How is a multiple choice question structured?
A multiple choice question comprises a question followed by a number - usually 4 or 5 - of which only one is correct and the others are not. The candidate has to select the correct one.
What is the probability of an incisional hernia after open-heart bypass surgery?
I have had two bypass surgeries. I didn't have an incisional hernia after the first one 8 years ago. But about 6 months after the second one the hernia began developing. It is large now, just under the breast bone, but the Doctors say it isn't large enough to do anything about it. It is about 4 inches across and bulges out about 1 inch in a ballooning manner. Wheeeee....
What is probabilty of drawing a ten from a deck of 52 playing cards?
7% chance
---------------------------------------------------------------------------------------------
There are 4 tens in a deck of 52 cards.
So the probability of drawing a ten from the deck is P(x=10) = 4/52 = 0.0769230...
P(x=10) ≈ 7.69%.
What is probability of getting more than 2 on one roll of a die?
Total number of possible outcomes = 6.
Number of them that are more than 2 = 4. (3, 4, 5, and 6)
Probability of rolling more than 2 = 4/6 = 2/3 = 662/3%
No. Drawing a red ball and then a blue ball from a bag are not mutually exclusive events, because the first event changes the probability of the second event, unless you return the first ball to the bag before drawing the second ball.
What is the probability of pulling out a slip of paper that has a vowel written on it?
The total number of slips of paper with vowels written on them
-------------------------------------------------------------------------------
The total number of slips of paper
This is assuming that the probability of pulling out each individual slip is the same for all slips.
How do you determine sample size given the standard error of 5?
There is not enough information to give an answer.
You need to know something about the distribution of the variable in the population. It is reasonable to assume that it is Gaussian (Normal)? Secondly, what power do you require of the test statistic? In other words, what level of significance do you require and that, in turn, depends on the "cost" of getting it wrong!
Reflexes are a primary function of the?
The primary function of reflexes is to prevent injury to your body and protect you from things that may be harmful to you and your body.
The basic formula we need is the Standard Normal Distribution formula (see related link) which is Z = (x - mean)/std dev. Solve for x, and we get: x = mean + z*std dev. From the Z table (see related link), area = .025 (your 2.5% in decimal form) we get the z value of -1.96. We can now solve for x. X = 400 + (-1.96)*50 or 302. Therefore, 302 is the point (value) below which 2.5% expenses fell (I assume it is $302).
What are the odds of winning millions in the lottery?
Very poor. New York Lotto, for instance, is 6 balls in a field of 54, with two games played for a dollar. The odds of hitting the jackpot is thus 2 in 25,827,165.
How many people have the birthday February.29?
You have to be a EXCELLANT matamatician to figure that out or travel the world and ask people but my guess would be about a few thousands or one million
Probability is 1/6 ~= .1667 = 16.67%
Since every month except May, June, July and August has a R in its name, the odds of someone born in a month whose name contains an R is roughly 2/3. The exact number is closer to 66.32% (242.2425 ÷ 365.2425).
What are the odds of getting five blacks in a row on the roulett table?
In American roulette there are 18 black pockets on the wheel, 18 red pockets, a "0" pocket and a "00" pocket. So there are 38 pockets in total. This makes the probability of getting a black on a single spin 18/38=0.474 (i.e. about 47.4%).
The probability of getting five blacks in a row is therefore (0.474)5=0.0238 (2.4%).
The odds (ratio of probable success to failure) of getting five blacks is therefore (1-0.0238)/0.238 or roughly 41 to 1 against.
Does empirical rules always apply to discrete probability distribution?
the empirical rules of probablility applies to the continuous probability distribution
What is the probability of drawing a red card and a spade in two draws without replacing?
The probability of drawing a red card followed by a spade is (1 in 2) times (1 in 4), or 1 in 8, or 0.125.
The probability of drawing a spade followed by a red card is (1 in 4) times (1 in 2), or 1 in 8, or 0.125.
Since you have two distinct desired outcomes, add them together, giving a probability of drawing a red card and a spade of 0.25.
You are tossing three coins and want to determine the probability for getting at least one head?
Pr(At least one head in three tosses) = 1 - Pr(No heads in three tosses)
= 1 - Pr(Three tails in three tosses) = 1 - (1/2)*(1/2)*(1/2) = 1 - 1/8 = 7/8 or 0.875 or 87.5%
Who knows I have never seen that place but
I can't say it's ugly that's mean!
How do you draw 42 cans of juice in an array OF MATH?
One example is making 7 cans of juice in each of six rows:
•••••••
•••••••
•••••••
•••••••
•••••••
•••••••
It is because the incumbent is removed for some reason or another.
Why is experimental probability called experimental probability?
Because it is the process of deriving probability through repeated experiments.
Probability of tossing a 5 in a number cube?
The probability of an event is the number of favourable outcomes divided by the total number of possible outcomes. What is the total number of possible outcomes of tossing a number cube? 6 How many outcomes are favourable to the event of getting a five? 1 So the prob is 1/6 or 0.16667
Can you put an ace spade with any card in black jack?
In Black Jack, the Ace of Spades is worth the same as any other Ace, 1 or 11, as needed. It has no special meaning. A "Black Jack" is any combination of two cards that total 21, so it is any Ace plus any face card or ten.